OpenStudy (anonymous):

Scalar multiplication is when you multiply two matrices of different dimensions. true or false?

OpenStudy (datanewb):

False. Multiplying two matrices is matrix multiplication. Scalar multiplication involves multiplying a vector or matrix by a scalar value, hence changing the scale of the vector. ie, \[4*\left[\begin{matrix}1 \\2\end{matrix}\right] = \left[\begin{matrix}4 \\8\end{matrix}\right]\] is an example of scalar multiplication where 4 is the scalar value multiplying the vector (1,2).